How many cubic yards of concrete will be used for 2500 feet of pavement that is 9 inches thick and 12 feet wide?
Answer by checkley77(12844)

I'll assume the 2500 ft. is the length.
2500*12*3/4=22,500 yd^3 of cement will be used.
1 cubic yd=3*3*3=27 ft^3
22,500*27=607,500 ft^3 of cement will be used.